ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 259, 2024

Alexandrino, M. M., B. Alves and M. A. Javaloyes On equifocal Finsler submanifolds and analytic maps

203

Alves, B. See M. M. Alexandrino Alves, C. O. and C. Ji Multi-peak positive solutions for a logarith-

mic Schrödinger equation via variational methods

835

Anderson, R. M., H. Duanmu and A. Smith Mixing times and hitting times for general Markov processes

759

Andruskiewitsch, N. On pointed Hopf algebras over nilpotent groups

169

Berele, A. and S. Catoiu Specialized symmetric functions and trace identities

503

Betsumiya, K., C. H. Lam and H. Shimakura Automorphism groups of cyclic orbifold vertex operator algebras associated with the Leech lattice and some non-prime isometries

621

Biagi, S. and A. Bonfiglioli Global heat kernels for parabolic homogeneous Hörmander operators

89

Bonfiglioli, A. See S. Biagi Borovoi, M. and D. A. Timashev Galois cohomology and compo-

nent group of a real reductive group

937

Boulanger, A. and S. Ghazouani SL_{2}(ℝ)-dynamics on the moduli space of one-holed dilation tori

1

Braga, B. M. Towards a theory of coarse geometry of operator spaces  527

Burguet, D. and R. Shi Mean dimension of continuous cellular automata

311

Catoiu, S. See A. Berele Chen, F., S. Tan and N. Yu Extended affine Lie algebras, vertex

algebras and equivariant ϕ-coordinated quasi-modules

347

De Medts, T. and J. Meulewaeter Inner ideals and structurable algebras: Moufang sets, triangles and hexagons

33

Duanmu, H. See R. M. Anderson Garbe, F., D. Král’ and A. Lamaison Hypergraphs with minimum

positive uniform Turán density

701

Garkusha, G. Group schemes and motivic spectra  727

Ghazouani, S. See A. Boulanger

Gut, F. and A. Joó Large vertex-flames in uncountable digraphs  401

Iglesias-Zemmour, P. Čech–De Rham bicomplex in diffeology  239

Javaloyes, M. A. See M. M. Alexandrino Ji, C. See C. O. Alves Joó, A. See F. Gut

Kastis, E. and S. C. Power The parabolic algebra revisited  559

Král’, D. See F. Garbe Kuca, B. Multidimensional polynomial Szemerédi theorem in finite

fields for polynomials of distinct degrees

589

Kundu, R. and A. Singh Generating functions for the powers

in GL(n, q) 887

Lahiri, A. and C. Sorensen Rigid vectors in p-adic principal series representations

427

Lamaison, A. See F. Garbe Lam, C. H. See K. Betsumiya Leite, E. and M. Montenegro Principal eigenvalues and eigenfunc-

tions to Lane–Emden systems on general bounded domains

277

Meulewaeter, J. See T. De Medts Montenegro, M. See E. Leite Power, S. C. See E. Kastis

Prokhorova, M. Spectral sections  651

Reid, C. D. Totally disconnected locally compact groups with just infinite locally normal subgroups

461

Shi, R. See D. Burguet Shimakura, H. See K. Betsumiya Singh, A. See R. Kundu Smith, A. See R. M. Anderson Sorensen, C. See A. Lahiri Tan, S. See F. Chen Timashev, D. A. See M. Borovoi Usuba, T. The list-chromatic number and the coloring number of un-

countable graphs

129

Yu, N. See F. Chen